5. If a new program could increase regular customers to of the total base and increase the average number of transac tions on Sunday, Monday, Wednesday, and Thursday to 125 what would it be worth (in revenues) to Video World?
Answer #1

Let the total base B be 100, original regular customers C1 be 10, original average number of transactions T1 be 100 and there be 7 days in a week.

Hence original revenue per day RD1 = C1 X T1 = 10 X 100 =1000

and original revenue per week would be RW1 = RD1 X Days in week = 1000 X 7 = 7000

As the new program increases the customers to 20 % of total base, hence current number of customers,

C2 = 20 % of B = 20% of 100 = 20

Current average number of transactions 4 days a week T2 = 125

Hence, current revenue per day RD2 = C2 X T2 = 20 X 125 = 2500

current revenue for such 4 days R4 = RD2 X 4 = 2500 X 4 = 10000

On the remaining 3 days, current revenue per day is like before 1000 and for 3 such days, it is R3 = 1000 X 3 = 3000

Hence, total current revenue for the week RW2 = R3 + R4 = 3000 + 10000 = 13000

For video world there would be worth a gain of RW2 - RW1 revenue, which means,

13000 - 7000 = 6000 units of revenue